# Robust Molecular Property Prediction via Densifying Scarce Labeled Data

### v3

@@ A.2 Implementation details of the ablation study (Table 2)
Since ablation studies (1) and (3) include the set function $\mu_{\lambda}$, we report the performance of each using both DeepSets (Zaheer et al., 2017) and Set Transformer (Lee et al., 2019) as implementations of $\mu_{\lambda}$.
@@
A.3 Hyperparameter tuning details
For hyperparameter settings, all methods (including ours, Mixup, Manifold Mixup, Mixup with outlier exposure, and Manifold Mixup with bilevel optimization) reported in Table 1 used a learning rate of $1\times 10^{-3}$ for $f_{\theta}$, and a learning rate of $1\times 10^{-5}$ for the mixers $\mu_{\lambda}$ and $\mu_{\alpha}^{\text{lin}}$, where the learning rate of $\mu_{\alpha}^{\text{lin}}$ is only used for Manifold Mixup (w/ bilevel optim.). The number of layers $(L)$ was set to 3, the hidden dimension $(H)$ to 64, the dropout rate to 0.5, and the optimizer to adamwschedulefree (Defazio et al., 2024). For methods utilizing bilevel optimization, we set the number of inner-loop iterations to 10 and outer-loop iterations to 50. We furthure perform hyperparameter search for all the models with the same hyperparameter search space as shown in Table 3. Since Mixup, Manifold Mixup do not utilize $\mathcal{D}_{\text{mvalid}}$, the hyperparameter search was conducted only over $M$.
@@ A.4 Computing resources
Appendix B t-SNE Visualization Details
B.1 Embedding generation procedure
@@
In Figure 2, we visualize the penultimate-layer embeddings $\mathbf{z}_{i}=x_{i}^{(L-1)}$ using t-SNE (Van der Maaten & Hinton, 2008), across three models that interpolate $\mathcal{D}_{\text{train}}$ with $\mathcal{D}_{\text{context}}$. For each trained model, we generate the following embeddings: $\mathbf{Z}_{\text{joint}}$: obtained by passing both $x_{i}$ and $C_{i}$ to the model; $\mathbf{Z}_{\text{input}}$: obtained by passing only $x_{i}$; $\mathbf{Z}_{\text{context}}$: obtained by passing only $C_{i}$. This allows us to observe how the model handles each type of input. To add an additional qualitative (OOD) evaluation, we utilize one of the Merck datasets (Kulkarni et al., 2012) that was not used during training (see Table 4). Using this OOD dataset, we generate the embeddings of $\mathbf{Z}_{\text{ood}}$, by only passing $x_{i}^{\text{ood}}$ to the model. In addition to Figure 2, we also provide extra examples in Figures 3, 4, 5, 6, 7 and 8.
B.2 t-SNE plots across all datasets
Here, we present the t-SNE plots for all datasets used in our experiments, as listed in Table 1. For each in-distribution model, we visualize the corresponding two out-of-distribution datasets. For our method, among the DeepSets (Zaheer et al., 2017) and Set Transformer (Lee et al., 2019) variants, we visualize the model that achieved the highest performance in Table 1.
@@
Appendix C Model Structure Overview
(a) Training and Test-time Framework
(b) Bilevel Optim. during Training

### v4

@@ B.2 Implementation details of the ablation study (Table 2)
Since ablation studies (1) and (3) include the set function $\mu_{\lambda}$, we report the performance of each using both DeepSets (deepsets) and Set Transformer (set-transformer) as implementations of $\mu_{\lambda}$.
@@
B.3 Hyperparameter tuning details
For hyperparameter settings, all methods (including ours, Mixup, Manifold Mixup, Mixup with outlier exposure, and Manifold Mixup with bilevel optimization) reported in Table 1 used a learning rate of $1\times 10^{-3}$ for $f_{\theta}$, and a learning rate of $1\times 10^{-5}$ for the mixers $\mu_{\lambda}$ and $\mu_{\alpha}^{\text{lin}}$, where the learning rate of $\mu_{\alpha}^{\text{lin}}$ is only used for Manifold Mixup (w/ bilevel optim.). The number of layers $(L)$ was set to 3, the hidden dimension $(H)$ to 64, the dropout rate to 0.5, and the optimizer to adamwschedulefree (adamsched). For methods utilizing bilevel optimization, we set the number of inner-loop iterations to 10 and outer-loop iterations to 50. We furthure perform hyperparameter search for all the models with the same hyperparameter search space as shown in Table 3. Since Mixup, Manifold Mixup do not utilize $\mathcal{D}_{\text{mvalid}}$, the hyperparameter search was conducted only over $M$.
@@ B.4 Computing resources
Appendix C t-SNE Visualization Details
C.1 Embedding generation procedure
@@
In Figure 2, we visualize the penultimate-layer embeddings $\mathbf{z}_{i}=x_{i}^{(L-1)}$ using t-SNE (tsne), across three models that interpolate $\mathcal{D}_{\text{train}}$ with $\mathcal{D}_{\text{context}}$. For each trained model, we generate the following embeddings: $\mathbf{Z}_{\text{joint}}$: obtained by passing both $x_{i}$ and $C_{i}$ to the model; $\mathbf{Z}_{\text{input}}$: obtained by passing only $x_{i}$; $\mathbf{Z}_{\text{context}}$: obtained by passing only $C_{i}$. This allows us to observe how the model handles each type of input. To add an additional qualitative (OOD) evaluation, we utilize one of the Merck datasets (merck_activity) that was not used during training (see Table 4). Using this OOD dataset, we generate the embeddings of $\mathbf{Z}_{\text{ood}}$, by only passing $x_{i}^{\text{ood}}$ to the model. In addition to Figure 2, we also provide extra examples in Figures 4, 5, 6, 7, 8 and 9.
C.2 t-SNE plots across all datasets
Here, we present the t-SNE plots for all datasets used in our experiments, as listed in Table 1. For each in-distribution model, we visualize the corresponding two out-of-distribution datasets. For our method, among the DeepSets (deepsets) and Set Transformer (set-transformer) variants, we visualize the model that achieved the highest performance in Table 1.